10. A continuous random variable X having values only between 0 and 4 has density function given by f (x) = 1/2 - ax ; where “a”

is constant. Calculate “a”. Find P(1<X<2)


1
Expert's answer
2021-12-07T13:41:29-0500

i.


"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{0}^{4}(\dfrac{1}{2}-ax)dx"

"=[\dfrac{x}{2}-\dfrac{ax^2}{2}]\begin{matrix}
 4 \\
 0
\end{matrix}=2-8a=1"

"a=\dfrac{1}{8}"

ii.


"P(1<X<2)=\displaystyle\int_{1}^{2}(\dfrac{1}{2}-\dfrac{1}{8}x)dx"

"=[\dfrac{x}{2}-\dfrac{x^2}{16}]\begin{matrix}
 2 \\
 1
\end{matrix}=1-\dfrac{1}{4}-(\dfrac{1}{2}-\dfrac{1}{16})"

"=\dfrac{5}{16}"